OpenStudy (anonymous):

@shivam_bhalla What is the initial concentration (in M) when 18 mL of a iron(III) chloride solution was diluted with 15.2 mL of solvent to a final concentration of 0.639 M? Assume the volumes are additive. -For the sake of a metal :) -

OpenStudy (anonymous):

Let x be the intial concentration no of milli moles = 18(x) Now new volume= 18+15.2 New concentration = no of milli moles / new volume WKT new concentration = 0.639 We get 18(x) / (18+15.2) = 0.639 18(x)=0.639 * 33.2 18(x)=21.215 x = 1.179 milli moles